The claim is that the proportion of peas with yellow pods is equal to 0.25​ (or 25%). The sample statistics from one experiment include 550 peas with 160 of them having yellow pods. Find the value of the test statistic.
Answer #1

Solution :

This hypothesis test is a two tailed test .  

The null and alternative hypothesis is

H0 : p = 0.25

Ha : p \neq 0.25

\hat p = x / n = 160 / 550 = 0.2909

1 - P0 = 1 - 0.25 = 0.75

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.2909 - 0.25 / [\sqrt(0.25 (0.75)) / 550]

= 2.216

The test statistic is 2.216